# Gateformer: Advancing Multivariate Time Series Forecasting through Temporal and Variate-Wise Attention with Gated Representations

### v2

@@ 4.1 FORECASTING RESULTS
In this section, we conduct extensive experiments to evaluate the forecasting performance of our proposed model together with advanced forecasters. We carefully choose nine well-acknowledged forecasting models as our benchmark, including (1) Transformer-based methods: Autoformer [Wu et al., 2021], FEDformer [Zhou et al., 2022], Stationary [Liu et al., 2022c], Crossformer [Zhang and Yan, 2023], PatchTST [Nie et al., 2023], iTransformer [Liu et al., 2024b]; (2) Linear-based methods: DLinear [Zeng et al., 2023]; and (3) TCN-based methods: SCINet [Liu et al., 2022a], TimesNet [Wu et al., 2023].
Experiment settings All models were evaluated under a consistent experimental setup, with a look-back window of $L=96$ and a fixed training duration of 10 epochs. The prediction lengths were $T\in\{3,6,12,24\}$ for the PEMS dataset; $T\in\{96,192,336,720\}$ for other datasets. Baseline results, except for the PEMS dataset, are from Liu et al. [2024b]. For the PEMS dataset, we specifically cite the results for prediction lengths $T\in\{12,24\}$ from Liu et al. [2024b]. All experiments were implemented in PyTorch [Paszke et al., 2019] and conducted on a single NVIDIA A100 40GB GPU. We reproduced the baseline models for comparison using the implementation from the TimesNet [Wu et al., 2023] repository.
Main results Table 2 presents multivariate forecasting results averaged across four prediction lengths, with best results highlighted in red and second-best in underlined. Lower MSE/MAE values indicate more accurate predictions. Our model consistently achieves state-of-the-art performance across all datasets, demonstrating its effectiveness in multivariate forecasting. Full results are provided in Appendix G. The comparison between Gateformer, iTransformer [Liu et al., 2024b], and Crossformer [Zhang and Yan, 2023] is noteworthy. Although iTransformer and Crossformer explicitly capture cross-variate dependencies, they underperform compared to our method. iTransformer focuses solely on capturing cross-variate correlations while neglecting cross-time dependency modeling, highlighting its limitation in encoding a variate’s time series through linear projection alone. Crossformer, similar to our method, divides each variable’s series into patches to model cross-time dependencies, followed by capturing cross-variate interactions across patches. However, its moderate performance suggests that modeling multivariate correlations at such a granular level may introduce unnecessary noise. In contrast, our method first encodes input series into variate-wise representations that integrate both cross-time dependencies and global temporal patterns. These representations are then used to model cross-variate dependencies. However, the flexibility to learn cross-variate relationships can be a double-edged sword, as it requires more training data to do so effectively [Nie et al., 2023, Han et al., 2024], leading to moderate performance on smaller datasets. While iTransformer and Crossformer struggle with low-dimensional datasets like ETT, our model performs on par with PatchTST [Nie et al., 2023], a representative channel-independent model that does not capture cross-variate dependencies. This strength can be attributed to our gating mechanism, which dynamically regulates the degree of cross-variate correlation, a capability that is especially critical in limited data scenarios. In summary, our approach excels at modeling both temporal and multivariate dependencies, making it highly effective for real-world time series forecasting tasks.
4.2 MODEL ANALYSIS
Ablation Study of Model Components To investigate the impact of each component in our method, we conducted ablation experiments on eight datasets: ETT (4 subsets), Weather, Traffic, Electricity, and Exchange. Specifically, we compared our model with three modified versions:
  • w/o Temporal-wise Attn. Removes the embeddings that encode temporal dependencies from the variate representation, leaving only embeddings that capture global temporal patterns.
  • w/o Gate in Variate-wise Attn. Removes the gating mechanism in cross-variate dependency modeling.
  • w/o Global Temporal Embeddings. Removes the embeddings that encode global temporal patterns, retaining only the embeddings that capture temporal dependencies in the variate representation.
As shown in Table 3, Gateformer (first row), utilizing both temporal-wise and channel-wise attention with gated representations, achieves the best overall performance. Notably, removing temporal-wise correlation (second row) leads to the poorest performance, underscoring its critical role. The primary performance advantage of our method over iTransformer stems from its ability to model temporal-wise dependencies. While removing the gating mechanism for cross-variate correlation (third row) improves performance slightly on high-dimensional datasets like Traffic and Electricity, where abundant training data allows effective learning of cross-dimensional dependencies, it significantly degrades performance on smaller datasets like ETT. Overall, our proposed method consistently delivers optimal results across both high-dimensional and low-dimensional datasets.
@@ 4.3 Framework Generalizability
Capturing cross-variate correlations is a key feature of our model. However, the quadratic complexity of self-attention becomes a computational bottleneck as the number of variates increases, a common challenge in real-world applications. Several prior works [Li et al., 2021, Wu et al., 2021, 2022, Zhou et al., 2022] have introduced advanced attention mechanisms to mitigate this issue. Since our model retains the standard Transformer architecture, it can seamlessly integrate these advanced attention modules from Transformer-based forecasters by substituting the conventional attention mechanism. We demonstrate this versatility by incorporating advanced attention mechanisms from models like Autoformer [Wu et al., 2021] and Flowformer [Wu et al., 2022], as well as extending our framework to LLM-based models such as GPT4TS [Zhou et al., 2023] to enhance their multivariate correlation modeling.
Experiments on the Weather and Electricity datasets (look-back window $L=96$, prediction lengths $T\in{96,192,336,720}$) show significant improvements over the original models: 19.9% on Autoformer, 20.7% on Flowformer, and 10.8% on GPT4TS, as shown in Table 4. These results demonstrate our framework’s effectiveness and adaptability across different model architectures. Detailed results are provided in Appendix D.
4.4 REPRESENTATION LEARNING
We evaluate the representations learned by our model, focusing on their generalization and transferability across datasets. The PEMS datasets are used as a benchmark, as they comprise four high-dimensional time series subsets, each containing a substantial amount of data to ensure robust and reliable evaluation results.
Transfer learning To assess representation transferability, we pre-trained the model on the largest dataset (PEMS07) for 10 epochs, then transferring it to PEMS03, PEMS04, and PEMS08 datasets. As shown in Table 5, the model’s zero-shot forecasting performance is comparable to most competitive baseline models. A key strength of our method lies in its ability to effectively capture cross-variate correlations, enabling superior generalization and transferability to unseen datasets. Furthermore, pre-training followed by fine-tuning for a few epochs consistently outperforms training solely on downstream datasets, demonstrating the value of transferable representations learned during pre-training.
4.5 MODEL EFFICIENCY
Memory-Efficient Training Strategy Our model captures multivariate correlations by applying the attention mechanism along the variate dimension. However, the quadratic complexity of the attention mechanism becomes a computational bottleneck when training models on high-dimensional datasets such as Traffic. To mitigate this, we implement a computationally efficient training strategy that randomly samples a subset of variates for each batch, allowing the model to train on only the selected variates. During inference, the model generates forecasts for all variates. This random sampling of variates acts as a form of regularization, enabling the model to learn robust and generalizable patterns. As shown in Figure 3, the forecasting performance remains stable across various sampling ratios, while the memory footprint is significantly reduced.
Model Efficiency To evaluate the efficiency of our model, we compare the forecasting accuracy, training time, and memory footprint of the following models: Gateformer, Gateformer with the memory-efficient training strategy, iTransformer [Liu et al., 2024b], PatchTST [Nie et al., 2023], Crossformer [Zhang and Yan, 2023], TimesNet [Wu et al., 2023], and DLinear [Zeng et al., 2023]. For a fair comparison, all models use the same hidden dimension and batch size on a representative high-dimensional dataset (Traffic).
Figure 4: Model efficiency comparison on Traffic dataset.
@@ 5 CONCLUSION AND FUTURE WORK
We propose a method that employs temporal and variate-wise attention modules to effectively capture both cross-time and cross-variate correlations. Gating mechanisms are incorporated to regulate the flow of information, enabling the model to focus on the most relevant features for accurate predictions. Experimentally, our approach achieves state-of-the-art performance and demonstrates flexibility, delivering up to a 20.7% performance improvement when integrated with other Transformer-based forecasters. In the future, we aim to explore large-scale pre-training and extend our method to a wider range of time series analysis tasks.
@@
All implementation details, including dataset descriptions, metrics, models, and experiment configurations, are provided in Appendix A. The code is submitted as Supplementary Material to the OpenReview system.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Experimental Details
A.1 Datasets
We assess long-term forecasting performance on ten widely used benchmarks, including the four ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Weather, Electricity, Traffic, Exchange (from Wu et al. [2021]), and Solar-Energy (from Lai et al. [2018]). Additionally, we evaluate short-term forecasting on the four PEMS datasets (PEMS03, PEMS04, PEMS07, PEMS08) as used in Liu et al. [2022a]. A brief description of each dataset is provided below.
  • ETT Dataset comprises two sub-datasets: ETTh and ETTm, which were collected from electricity transformers. Data were recorded at 15-minute and 1-hour intervals for ETTm and ETTh, respectively, spanning from July 2016 to July 2018.
  • Weather Dataset includes 21 meteorological indicators recorded every 10 minutes during 2020.
  • Electricity Dataset contains hourly electricity consumption data for 321 customers from 2012 to 2014.
  • Traffic Dataset provides hourly road occupancy rates measured by various sensors on San Francisco Bay Area freeways, collected by the California Department of Transportation.
  • Exchange Dataset collects daily exchange rates from 8 countries, spanning from 1990 to 2016.
  • Solar-Energy records the solar power production of 137 PV plants in 2006, which are sampled every 10 minutes.
  • PEMS Dataset contains public traffic network data from California, recorded at 5-minute intervals.
@@ A.1 Datasets
We adopt the same data processing and train-validation-test split protocol as TimesNet [Wu et al., 2023], ensuring datasets are strictly divided chronologically to prevent data leakage. For forecasting, the lookback series length is fixed at $L=96$ across all datasets. Prediction lengths are set as follows: $T\in\{96,192,336,720\}$ for ETT, Weather, ECL, Traffic, Solar-Energy, and Exchange; $T\in\{3,6,12,24\}$ for PEMS. Detailed dataset information is provided in Table 6.
@@
A.2 Implementation Details
All experiments were implemented in PyTorch [Paszke et al., 2019] and conducted on a single NVIDIA A100 40GB GPU. We used MSE as the loss function, a batch size of 8, and the Adam optimizer. The initial learning rate was selected from $\{10^{-4},5\times 10^{-4},10^{-3}\}$, with training running for a maximum of 10 epochs and early stopping applied if validation loss did not improve within three epochs. The number of encoder blocks was chosen from $L\in\{1,2,3,4\}$, and the hidden state dimension was selected from $\{64,128,256,512,1024\}$. Evaluation metrics included mean square error (MSE) and mean absolute error (MAE), with results averaged over three random seeds. Baseline models were reproduced using the TimesNet [Wu et al., 2023] repository, following configurations from their respective original papers or official code.
@@ A.2 Implementation Details
Appendix B Varying Look-back Window
In principle, expanding the look-back window increases the receptive field, which should enhance forecasting accuracy. However, previous studies [Zeng et al., 2023] have shown that most Transformer-based models fail to demonstrate this expected improvement, revealing their limitations in processing extended temporal sequences. Our model, in contrast, effectively utilizes the increased receptive field, achieving lower MSE scores with longer look-back windows, as demonstrated in Figure 5.
Appendix C Hyperparameter Sensitivity
@@
Appendix D Framework Generalizability
As discussed in Section 4.3, our proposed framework can be seamlessly integrated into other Transformer-based and LLM-based models to consistently improve their performance, with full forecasting results shown in Table 7. Since our model explicitly captures cross-variate dependencies, the variate-wise attention becomes a computational bottleneck when training on high-dimensional datasets such as Electricity, Traffic, and Solar-Energy. To mitigate this, we replace the conventional quadratic-time attention mechanism with the linear-time attention proposed in Flowformer [Wu et al., 2022] and compare its performance against competitive baselines. These results are provided in Figure 7.
Models
Autoformer
Flowformer
GPT4TS
[2021]
[2022]
[2023]
Metric
MSE
MAE
MSE
MAE
MSE
MAE
Electricity Original 96 0.201 0.317 0.215 0.320 0.197 0.290
192 0.222 0.334 0.259 0.355 0.201 0.292
336 0.231 0.338 0.296 0.383 0.217 0.309
720 0.254 0.361 0.296 0.380 0.253 0.339
Avg
0.227
0.338
0.267
0.359
0.217
0.308
+ Our framework 96 0.173 0.259 0.144 0.237 0.150 0.241
192 0.180 0.267 0.160 0.251 0.163 0.254
336 0.198 0.285 0.175 0.269 0.180 0.272
720 0.241 0.320 0.205 0.296 0.214 0.301
Avg
0.198
0.283
0.187
0.274
0.177
0.267
Weather Original 96 0.266 0.336 0.182 0.233 0.203 0.244
192 0.307 0.367 0.250 0.288 0.247 0.277
336 0.359 0.395 0.309 0.329 0.297 0.311
720 0.419 0.428 0.404 0.385 0.368 0.356
Avg
0.338
0.382
0.286
0.308
0.279
0.297
+ Our framework 96 0.175 0.218 0.171 0.213 0.175 0.214
192 0.219 0.258 0.220 0.258 0.228 0.260
336 0.281 0.300 0.275 0.297 0.282 0.301
720 0.353 0.350 0.353 0.349 0.359 0.352
Avg
0.257
0.281
0.255
0.279
0.261
0.282
@@ Appendix D Framework Generalizability
Appendix E Unified Cross-Dataset Training
@@
We selected the PEMS dataset as our benchmark as it contains four high-dimensional time-series subsets (PEMS03, PEMS04, PEMS07, and PEMS08), each with substantial data that ensures robust evaluation. Our model includes temporal-wise and variate-wise self-attention modules. The temporal-wise attention module captures intra-series dependencies for each series independently. The variate-wise attention module operates on the variate dimension, enabling unified training across datasets with varying numbers of variates. To investigate the benefits of data scaling, we trained our model on four progressively larger combined datasets: (1) PEMS08, (2) PEMS08 + PEMS04, (3) PEMS08 + PEMS04 + PEMS03, and (4) PEMS08 + PEMS04 + PEMS03 + PEMS07. These datasets were combined in ascending order of dimensionality. As shown in Figure 8, forecasting accuracy improved consistently across all PEMS datasets with larger training set sizes. This highlights our model’s ability to learn generalizable and transferable representations through unified training while preventing catastrophic forgetting, a common challenge in combined training Liu et al. [2024a]. Our results highlight the potential of building large-scale multivariate forecasting models trained across datasets using our framework.

### v3

@@ 3.1 Forecasting Results
We compare our model against nine SOTA methods: Transformer-based (Autoformer, FEDformer, Stationary, Crossformer, PatchTST, iTransformer), linear-based (DLinear), and TCN-based (SCINet, TimesNet).
Experimental Setup
All models use a look-back window of $L=96$ and are trained for 10 epochs. Prediction lengths are $T\in\{3,6,12,24\}$ for PEMS, and $T\in\{96,192,336,720\}$ for others. Baseline results follow Liu et al. (2024b).
Main Results
Table 1 shows that Gateformer consistently outperforms all baselines across 13 datasets. Although iTransformer and Crossformer capture cross-variate dependencies, they fall short due to limitations in modeling temporal dynamics or introducing noise through patch-wise interactions. Our method encodes both temporal and global patterns via gated representations, enabling robust cross-variate modeling. This flexibility allows strong performance across dataset scales. On low-dimensional datasets (e.g., ETT), Gateformer matches or exceeds PatchTST, a strong channel-independent baseline, thanks to its gated control of cross-variate interactions.
3.2 Ablation Study
We assess component contributions on eight datasets (ETT, Weather, Traffic, Electricity, Exchange) by removing: (1) Temporal-wise Attention (keeps only global patterns), (2) Gate in Variate-wise Attention, and (3) Global Temporal Embeddings (keeps only local patterns). Table 2 shows full Gateformer achieves the best results. Removing temporal attention significantly degrades performance, especially on small datasets, confirming its importance. Removing the variate-wise gate slightly improves performance on large datasets but harms it on smaller ones. Overall, the gated integration of temporal and variate-wise dependencies ensures strong, consistent forecasting performance.
@@ 3.3 Framework Generalizability
While capturing cross-variate dependencies is essential, the quadratic cost of self-attention limits scalability with many variates. Our Transformer-based framework addresses this by easily integrating efficient attention modules from models like Autoformer (Wu et al., 2021) and Flowformer (Wu et al., 2022), and by extending to LLM-based models like GPT4TS (Zhou et al., 2023) or foundation models such as Moment (Goswami et al., 2024). On Weather and Electricity datasets, our framework improves Autoformer by 19.9%, Flowformer by 20.7%, GPT4TS by 10.8% and Moment by 6.9%, as shown in Table 3. These gains highlight the flexibility and effectiveness of our approach.
@@ 4 Conclusion and Future Work
We present a model that captures both temporal and cross-variate dependencies using attention and gating mechanisms for improved forecasting. It achieves state-of-the-art results and boosts performance of existing Transformer-based models by up to 20.7%. Future work includes large-scale pre-training and broader time series applications.
@@
Code, dataset details, metrics, and experimental settings are included in Appendix B.
Appendix A Related work
Point-wise Representations with Cross-time Attention
Many prior works embed multiple variates at the same timestamp into point tokens and apply attention mechanisms to capture temporal dependencies among them. Examples include Autoformer (Wu et al., 2021), FEDformer (Zhou et al., 2022), and Pyraformer (Liu et al., 2022b), which focus on optimizing the quadratic complexity of self-attention. Autoformer detects sub-series similarities with $O(L\log L)$ complexity, FEDformer leverages frequency-domain sparsity with $O(L)$ complexity, and Pyraformer uses pyramidal attention for short- and long-term dependencies with $O(L)$ complexity. However, these models often fail in multivariate forecasting due to the lack of explicit semantic meaning in point tokens and the loss of cross-variate correlations when merging all variables into single temporal tokens.
Patch-wise Representations with Cross-time Attention PatchTST (Nie et al., 2023) addresses the lack of local semantic information through patching. It divides time series into patches to expand the receptive field and applies self-attention to capture cross-time dependencies among patches. PatchTST processes each variate independently, allowing unique attention patterns to be learned for each series. This approach has consistently improved performance on benchmarks and inspired recent large-scale time series models (Zhou et al., 2023; Das et al., 2024; Jin et al., 2024; Liu et al., 2024c). However, it focuses primarily on cross-time dependencies while neglecting critical cross-variate interactions.
Variate-wise Representations with Cross-variate Attention A notable example is iTransformer (Liu et al., 2024b), which expands the receptive field as an extreme form of patching. iTransformer encodes each variate’s time series into coarse variate-wise representations through linear projection and explicitly models cross-variate correlations among these representations. However, it does not account for cross-time (intra-series) dependencies.
Patch-wise Representations with Cross-time and Cross-variate Attention Crossformer (Zhang & Yan, 2023), a representative model, segments each variate’s series into a sequence of patches and captures both cross-time (intra-series) and cross-variate (inter-series) dependencies among them. However, modeling multivariate correlations at such a granular level can introduce unnecessary noise, potentially resulting in suboptimal forecasting performance.
Unlike previous works, our approach captures both cross-time and cross-variate correlations through variate-wise representations. To enhance the quality of these representations, we integrate gating mechanisms to regulate information flow, allowing the model to focus on the most relevant features for accurate predictions.
Appendix B Experimental Details
B.1 Datasets
@@
We assess long-term forecasting performance on ten widely used benchmarks, including the four ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Weather, Electricity, Traffic, Exchange (from Wu et al. (2021)), and Solar-Energy (from Lai et al. (2018)). Additionally, we evaluate short-term forecasting on the four PEMS datasets (PEMS03, PEMS04, PEMS07, PEMS08) as used in Liu et al. (2022a). We adopt the same data processing and train-validation-test split protocol as TimesNet (Wu et al., 2023), ensuring datasets are strictly divided chronologically to prevent data leakage. For forecasting, the lookback series length is fixed at $L=96$ across all datasets. Prediction lengths are set as follows: $T\in\{96,192,336,720\}$ for ETT, Weather, ECL, Traffic, Solar-Energy, and Exchange; $T\in\{3,6,12,24\}$ for PEMS. Detailed dataset information is provided in Table 4.
@@
B.2 Implementation Details
All experiments were implemented in PyTorch (Paszke et al., 2019) and conducted on a single NVIDIA A100 40GB GPU. We used MSE as the loss function, a batch size of 8, and the Adam optimizer. The initial learning rate was selected from $\{10^{-4},5\times 10^{-4},10^{-3}\}$, with training running for a maximum of 10 epochs and early stopping applied if validation loss did not improve within three epochs. The number of encoder blocks was chosen from $L\in\{1,2,3,4\}$, and the hidden state dimension was selected from $\{64,128,256,512,1024\}$. Evaluation metrics included mean square error (MSE) and mean absolute error (MAE), with results averaged over three random seeds. Baseline models were reproduced using the TimesNet (Wu et al., 2023) repository, following configurations from their respective original papers or official code.
@@ B.2 Implementation Details
Appendix C Varying Look-back Window
In principle, expanding the look-back window increases the receptive field, which should enhance forecasting accuracy. However, previous studies (Zeng et al., 2023) have shown that most Transformer-based models fail to demonstrate this expected improvement, revealing their limitations in processing extended temporal sequences. Our model, in contrast, effectively utilizes the increased receptive field, achieving lower MSE scores with longer look-back windows, as demonstrated in Figure 3.
Appendix D Hyperparameter Sensitivity
@@
Appendix E Framework Generalizability
As discussed in Section 3.3, our proposed framework can be seamlessly integrated into other Transformer-based and LLM-based models to consistently improve their performance, with full forecasting results shown in Table 3. Since our model explicitly captures cross-variate dependencies, the variate-wise attention becomes a computational bottleneck when training on high-dimensional datasets such as Electricity, Traffic, and Solar-Energy. To mitigate this, we replace the conventional quadratic-time attention mechanism with the linear-time attention proposed in Flowformer (Wu et al., 2022) and compare its performance against competitive baselines. These results are provided in Figure 5.
@@ Appendix E Framework Generalizability
Appendix F Unified Cross-Dataset Training
@@
We selected the PEMS dataset as our benchmark as it contains four high-dimensional time-series subsets (PEMS03, PEMS04, PEMS07, and PEMS08), each with substantial data that ensures robust evaluation. Our model includes temporal-wise and variate-wise self-attention modules. The temporal-wise attention module captures intra-series dependencies for each series independently. The variate-wise attention module operates on the variate dimension, enabling unified training across datasets with varying numbers of variates. To investigate the benefits of data scaling, we trained our model on four progressively larger combined datasets: (1) PEMS08, (2) PEMS08 + PEMS04, (3) PEMS08 + PEMS04 + PEMS03, and (4) PEMS08 + PEMS04 + PEMS03 + PEMS07. These datasets were combined in ascending order of dimensionality. As shown in Figure 6, forecasting accuracy improved consistently across all PEMS datasets with larger training set sizes. This highlights our model’s ability to learn generalizable and transferable representations through unified training while preventing catastrophic forgetting, a common challenge in combined training (Liu et al., 2024a). Our results highlight the potential of building large-scale multivariate forecasting models trained across datasets using our framework.
Appendix G REPRESENTATION LEARNING
We evaluate the representations learned by our model, focusing on their generalization and transferability across datasets. The PEMS datasets are used as a benchmark, as they comprise four high-dimensional time series subsets, each containing a substantial amount of data to ensure robust and reliable evaluation results. To assess representation transferability, we pre-trained the model on the largest dataset (PEMS07) for 10 epochs, then transferring it to PEMS03, PEMS04, and PEMS08 datasets. As shown in Table 5, the model’s zero-shot forecasting performance is comparable to most competitive baseline models. A key strength of our method lies in its ability to effectively capture cross-variate correlations, enabling superior generalization and transferability to unseen datasets. Furthermore, pre-training followed by fine-tuning for a few epochs consistently outperforms training solely on downstream datasets, demonstrating the value of transferable representations learned during pre-training.
Models Gateformer Models Trained from Scratch on Target Dataset
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Appendix H MODEL EFFICIENCY
H.1 Memory-Efficient Training Strategy
Our model captures multivariate correlations by applying the attention mechanism along the variate dimension. However, the quadratic complexity of the attention mechanism becomes a computational bottleneck when training models on high-dimensional datasets such as Traffic. To mitigate this, we implement a computationally efficient training strategy that randomly samples a subset of variates for each batch, allowing the model to train on only the selected variates. During inference, the model generates forecasts for all variates. This random sampling of variates acts as a form of regularization, enabling the model to learn robust and generalizable patterns. As shown in Figure 7, the forecasting performance remains stable across various sampling ratios, while the memory footprint is significantly reduced.
H.2 Model Efficiency
To evaluate the efficiency of our model, we compare the forecasting accuracy, training time, and memory footprint of the following models: Gateformer, Gateformer with the memory-efficient training strategy, iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), Crossformer (Zhang & Yan, 2023), TimesNet (Wu et al., 2023), and DLinear (Zeng et al., 2023). For a fair comparison, all models use the same hidden dimension and batch size on a representative high-dimensional dataset (Traffic).
Figure 8: Model efficiency comparison on Traffic dataset.
Appendix I Visualization